# AgenticTagger: Structured Item Representation for Recommendation with LLM Agents

### v1

@@ 4.2. Vocabulary Assignment
Once the feature vocabulary $A_{all}$ is built, the vocabulary assignment stage for AgenticTagger is straightforward, where we simply constrain the descriptors output by the LLMs to fall into the vocabulary by prompting, i.e., $g(X_{i})=\{a\in A\mid\pi(X_{i},A)\to a\}$ for any item raw attribute $X_{i}$. For example, in AgenticTagger, for the sake of implementation simplicity, we opt to reuse the same operation from building the vocabulary, $\text{ParallelAssign}(I,A_{all},\pi_{\text{Annotator}})$, where an LLM is prompted to assign features to an item by observing all features in its context, exploiting modern LLMs’ long-context and reasoning abilities (Zhao et al., 2025). However, depending on the existing LLM serving infrastructure, this step can also be implemented as a post-processing step, where we match an LLM generation to its nearest neighbor in the entire feature vocabulary $A_{all}$, if their similarity exceeds some threshold metric (e.g., edit distance). We provide further discussion of these design choices in Section 6.1, and show the advantage of the prompting-based method, despite the longer prompts at the assignment stage.
We note that due to the prompting nature of the vocabulary assignment stage, AgenticTagger also naturally supports arbitrary natural language instructions, offering controllability from system deployers’ side. As we later showcase in our experiments (Section 5), for example, when the vocabulary itself contains hierarchical structure, one could simply instruct LLM to select one best descriptor per taxonomy level for each item, where the produced vocabulary becomes a specialized semantic ID (Rajput et al., 2023; Hou et al., 2023, 2025b, 2025a; Singh et al., 2024), where each ID is a natural language descriptor.
@@ 4.3. Complexity Analysis
In AgenticTagger’s automated vocabulary building algorithm, the parallel calls to the annotator LLMs during the vocabulary creation phase dominate. In its most basic version, the algorithm will incur $O(C_{max}\cdot N\cdot b^{D_{max}})$ LLM calls, assuming an item will be assigned to an average of $b$ features at each depth level $d\in\{1,\dots,D_{max}\}$, which grows quickly as the depth of the feature hierarchy grows. However, the expressiveness of the feature hierarchy also grows exponentially w.r.t. the depth of the hierarchy, similar to other hierarchical discrete item representation methods (Rajput et al., 2023; Zhang et al., 2024c), and thus a few layers are often sufficient, as we later show in our experiments, 3-6 layers of features suffice good performance across public benchmarks (Table 2).
Further, since the goal of vocabulary building is solely to generate a set of feature descriptors, we can also heuristically constrain the branching factor by limiting how many features an item is assigned to at each layer, such as randomly assigning an item to one of its associated feature during the branching operation. In practice, we observe that a branching factor of 1 is also sufficient for good performance (Section 7.2), in which case the complexity w.r.t. LLM calls becomes $O(C_{max}\cdot N\cdot D_{max})$. Finally, compared to prior LLM-based document categorization methods (Wan et al., 2024; Pham et al., 2024), since AgenticTagger does not have sequential dependency between most of its dominating LLM operations (specifically, the $\pi_{\text{Annotator}}$ annotations), the overall execution time can be further reduced approximately linearly w.r.t. the number of concurrent threads in the ParallelAssign operation.
Table 2. AgenticTagger for Generative Recommendation on public benchmarks. For each metric, the best result is in bold and the second-best is underlined. Our model’s performance is reported across 3 runs. We provide additional comparison against crossing-based IDs and free-form generations in Figure 3 and Section 7.3.
@@
Since once built, the assignment stage simply involves looping through all items of interest, the assignment loop induces $O(|\mathcal{U}|)$ LLM calls, i.e., the cost of querying $\pi_{Annotator}$ in a single-pass manner over the item corpus. To this end, once the vocabulary is built, AgenticTagger can easily scale to large item corpora with parallelized calls to LLMs. Further, since the only difference between each LLM call during the assignment stage is the item description, the inference can, in principle, be further accelerated by infrastructure optimizations such as KV-Cache (Pope et al., 2023), which we leave for future work.
@@ 4.4. Modeling AgenticTagger Features
Once the process is complete, AgenticTagger assigns a fixed maximum number of descriptors to an item, where each “slot” represents a categorical feature. In this work, we opt to assume the best potential value for each slot is nominal, i.e., mutually exclusive, although there could be more than one suitable value at the same slot for each item (e.g., a song that is both “Rock” and “Blues”). These assigned descriptors also have an inherent order, where a preceding descriptor represents the broader, parent category of its child descriptors. For example, for a item to fit “Glam Rock” descriptor, it’s preceding, broader feature would be “Rock”. As we later show, this formulation allows these features to be natively used as item IDs such as semantic IDs and ID-like ranking features Section 5, as well as regular natural language term-based descriptors in use cases like bag-of-terms retrieval Section 6.1.
@@ 5. Main Experiments
In this section, we showcase the performance of the AgenticTagger-produced descriptor features by deploying them as discrete ID features in ML models, in particular, generative recommendation (Section 5.1) and ranking (Section 5.2). Meanwhile, we defer discussion on term-based retrieval (Section 6.1) and interpretability (Section 6.2) to later sections, alongside other analysis-focused research questions.
@@ 5.1. Generative Recommendation
We conduct our experiments across three domains from the Amazon Reviews dataset (McAuley et al., 2015): “Sports and Outdoors” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs), the dataset statistics are shown in Table 6. We adopt the commonly used “last-out” splitting (Kang and McAuley, 2018; Rajput et al., 2023; Hou et al., 2025a), where we hold out the last chronologically interacted item for each user as the test set, and the second-last item as the validation set. Following recent work (Rajput et al., 2023; Hou et al., 2023; Wang et al., 2024a), we evaluate the performance of models with Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. To eliminate randomness in model optimization, we repeat our experiments with 3 random seeds and report the average metrics. We use Gemini-2.5-flash and Gemini-2.5-flash-lite for the architect and annotator LLMs, respectively, in our main experiment, and discuss results from Gemini-1.5-flash and Gemini-1.5-flash-lite (specifically, the 001 series) where we achieved similarly competitive performance in Section B.1.
@@
In our main experiment (Table 2), we compare our method with a few streams of models: models with item IDs only, such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019); models enhanced by embedded features: FDSA (Zhang et al., 2019), $\text{S}^{3}$-Rec (Zhou et al., 2020); and models with discrete features (i.e., semantic IDs) derived from semantic information or collaborative information: VQ-Rec (Hou et al., 2023), TIGER (Rajput et al., 2023), HSTU (Zhai et al., 2024), and LLM-Indexer (Jin et al., 2024). We additionally compare our method with models that enhances raw semantic IDs via feature crossing, namely SPM-CID (Singh et al., 2024) and ActionPiece (Hou et al., 2025b) in Figure 3.
@@
Interestingly, the performance improvements are more significant in the Sports and CDs domain, with 8.9%-11.6% improvements against the closest competing model in NDCG@10, where as the improvement is moderate (2%) in the Beauty domain. We hypothesize that this could be due to AgenticTagger yields the most benefit when it is able to construct its vocabulary from a larger items corpus, where the natural language feedback aggregated via the Distill operation provides more stable signal. We further discuss this hypothesis in Section 7.2, where we show that AgenticTagger benefit from larger dataset with diverse items, even when doing so at the cost of pruning the vocabulary building process.
@@
In light of recent advances on showing the benefit of incorporating dataset collaborative filtering statistics directly into the tokenization process via feature crossing (Hou et al., 2025b; Singh et al., 2024), we also compare our method with two recent tokenization methods that merge frequently co-occurring semantic IDs into new IDs based on observed user behavior sequences, SentencePiece (Singh et al., 2024) and ActionPiece (Hou et al., 2025b). To be specific, these methods represent an alternative path towards better features for recommender systems: automated feature crossing based on user behavior. As shown in Figure 3, AgenticTagger remains a strong tokenization method, demonstrating the effectiveness of leveraging LLMs’ content understanding ability in tokenization in addition to collaborative filtering signals. Meanwhile, we note that AgenticTagger is orthogonal to these crossing-based techniques, and leaves crossing AgenticTagger features to future study.
@@ 5.2. Ranking
R@5 0.0833 $\pm$ 0.0068 0.1205 $\pm$ 0.0080
N@5 0.0589 $\pm$ 0.0051 0.0853 $\pm$ 0.0060
R@10 0.1163 $\pm$ 0.0079 0.1751 $\pm$ 0.0093
N@10 0.0695 $\pm$ 0.0052 0.1028 $\pm$ 0.0061
Besides generative recommendation, another frequent evaluation for recent discrete item ID generation frameworks is ranking (Singh et al., 2024; Zhu et al., 2021). To this end, we evaluate the effect of AgenticTagger-features on a private dataset, since publicly available ranking datasets generally anonymize item features (Zhu et al., 2025b), making AgenticTagger inapplicable. Similar to the generative recommendation experiments Section 5.1, we use the trained ranking model to perform click-through-rate-based ranking, and report performance in Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. For dataset construction, we collect interaction logs from a news-feed serving platform over a random subset of 1791 users, 15,830 items, and 47,260 interactions. We report our result as in Table 3, as shown, AgenticTagger brings significant improvement to ranking accuracy compared to raw item IDs, showcasing the effectiveness of our proposed approach in effectively compressing item semantic information.
6. Further Applications
6.1. Term-based Retrieval
Figure 4. Comparison between AgenticTagger (AT) and free-form generated discriptor-tags (FT) variants on term-based retrieval. Top: Recall@k results across baselines. Bottom: Performance-utilization trade-off. AgenticTagger achieves the best cardinality-performance trade-off, with further gain when combined with FT. Our method achieves up to 81.9% feature utilization compared to only 6.1% for free-form generation.
In term-based retrieval, user histories are aggregated by summing the bag-of-terms representation of each interacted item, and target documents are ranked by their dot product scores w.r.t. the user-history representation. We choose this setting since it is a model-free prediction pipeline that naturally handles large cardinalities well, allowing us to study the effect of methods with dramatically varying vocabulary sizes while eliminating randomness in model optimization.
We consider 3 layers of AgenticTagger features (AT), free-form generated descriptor tags (FT), and variants of FT that reduces its vocabulary by grouping semantically similar features via k-means clustering (FT+KM), and finally, combining AT and FT feateures via feature-crossing (AT+FT). For the AT+FT method specifically, we prompt the LLM to generate three free-form keywords in addition to the AgenticTagger vocabulary during the vocabulary assignment stage, to study the potential headroom for combining different methods. The free-form keywords are then crossed with level-2 descriptors in AgenticTagger to form additional features, for effective feature enhancement under moderate vocabulary size overhead. Experiments are conducted on a random subset user interaction logs over a week from the same subset in Section 5.2 that consists of 1791 users, 15,830 items, and 47,260 interactions after removing repeated item consumptions.
The results are shown in Figure 4. By comparing FT+KM and FT, we can see that a larger vocabulary benefits retrieval performance, showing the effectiveness of LLM-generated features. Interestingly, AT can already match the performance of FT, especially on coarse retrieval (R@50), showcasing its advantage in cardinality-sensitive scenarios. Finally, AT+FT achieves the best performance, despite having lower cardinality than vanilla FT (since only three keywords are generated per item), showcasing the potential value of combining AgenticTagger-features with other features such as keywords.
As in prior works for (numerical) discrete item representations for recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; Rajput et al., 2023; He et al., 2025), another important property to study for a given vocabulary codebook and its assignment is its vocabulary utilization rate, i.e., how many features are effectively used across the item corpus. While AgenticTagger, by its nature, guarantees that each feature it generates is assigned to at least one item, we believe it is meaningful in this case to examine the number of descriptor features that appear in more than one item, as only these features enable effective cross-item generalization. To this end, we report the vocabulary utilization rate and performance trade-off in the lower sub-figure of Figure 4. Specifically, both AT and AT+FT achieve a better performance-utilization trade-off, demonstrating that AgenticTagger’s design effectively improves the framework beyond free-form generation.
@@ 6.1. Term-based Retrieval
6.2. Interpretability and Controllability
As discussed earlier, a core advantage of LLM-generated features is their explainability, which naturally suits user-facing use cases, such as situations where one would like to control or constrain the matching mechanism of a recommender system. In this section, we discuss the value AgenticTagger features add to the interpretability and controllability of a recommender system via single-turn critique setting, where a user gives feedback to a recommender system, providing information on the item they would like to interact with next. In particular, this setting is a fundamental ability in many conversational recommender systems (Li et al., 2018; Wang et al., 2022; He et al., 2023; Zhu et al., 2025c), and is commonly manifested in production recommender systems as feedback with keyphrase-based rationales.
Inspired by recent work on using LLMs for user simulation in recommender systems (Bougie and Watanabe, 2025; Park, 2025; Zhang et al., 2025a; Wang et al., 2024e) and conversational recommendation (Wang et al., 2023; Zhu et al., 2025a), we opt to make use of an LLM as user simulator to evaluate whether AgenticTagger can enable a user to perform effective control over recommender system behavior. Specifically, we adopt a communication game with an information bottleneck setting, where an LLM (gemini-2.5-flash in practice) observes the descriptions of target items in its prompt, and is asked to select all potentially relevant top-level (i.e., level 1) features in the vocabulary, which generally consists of a small amount of keywords (e.g., 16 for the CDs and Vinyl domain, the largest public benchmark we evaluate on). We then generate the next item from a pre-trained generative recommender backbone using beam search, following standard practice, except we constrain its search space to be within the allowed level 1 features provided by the user simulator. While this interaction paradigm essentially “leaks” ground-truth information to the recommender, it verifies that (1) AgenticTagger produces a vocabulary to precisely identify meaningful subset of items, and (2) the produced vocabulary aligns well with the semantic understanding of an LLM, which are known to align well with human perception (Chang and Bergen, 2024), demonstrating the interpretability and controllability potential of AgenticTagger features.
As shown in Table 4, this constrained recommendation setting increases model performance significantly across datasets, showcasing that AgenticTagger features effectively communicate the nuances of items in the dataset, and importantly, in an interpretable manner, such that the behavior of the corresponding recommender system using AgenticTagger-features can be straightforwardly controlled by natural-language-bridged critiques.

### v2

@@ 4.2. Vocabulary Assignment
Once the feature vocabulary $A_{all}$ is built, the vocabulary assignment stage for AgenticTagger is straightforward, where we simply constrain the descriptors output by the LLMs to fall into the vocabulary by prompting, i.e., $g(X_{i})=\{a\in A\mid\pi(X_{i},A)\to a\}$ for any raw item attribute $X_{i}$. For example, in AgenticTagger, for the sake of implementation simplicity, we opt to reuse the same operation from building the vocabulary, $\text{ParallelAssign}(I,A_{all},\pi_{\text{Annotator}})$, where an LLM is prompted to assign features to an item by observing all features in its context, exploiting modern LLMs’ long-context and reasoning abilities (Zhao et al., 2025). As we later show in Section 5, the generated descriptors are indeed of sufficient quality to support good performance in downstream recommendation tasks. We provide further analysis and qualitative examples in Section B.4.
We note that in principle, there might be multiple descriptors of fit at each depth of the vocabulary hierarchy (e.g., a CD that is both “Rock” and “Blues”). To this end, we simply instruct the LLM to select a single best descriptor per level, enforcing structural constraints while preserving item information. As we later showcase in our experiments (Section 5), this is sufficient for good downstream recommendation performance, and we leave better collision-resolving methods to future work.
@@ 4.3. Complexity Analysis
In AgenticTagger’s automated vocabulary-building algorithm, the parallel calls to the annotator LLMs during the vocabulary creation phase dominate. In its most basic version, the algorithm will incur $O(C_{max}\cdot N\cdot b^{D_{max}})$ LLM calls, assuming an item will be assigned to an average of $b$ features at each depth level $d\in\{1,\dots,D_{max}\}$, which grows quickly as the depth of the feature hierarchy grows. However, the expressiveness of the feature hierarchy also grows exponentially w.r.t. the depth of the hierarchy, similar to other hierarchical discrete item representation methods (Rajput et al., 2023; Zhang et al., 2024c), and thus a few layers are often sufficient; as we later show in our experiments, 3-6 layers of features are sufficient for good performance across public benchmarks (Table 1).
Further, since the goal of vocabulary building is solely to generate a set of feature descriptors, we can also heuristically constrain the branching factor by limiting how many features an item is assigned to at each layer, such as randomly assigning an item to one of its associated features during the branching operation. In practice, we observe that a branching factor of 1 is also sufficient for good performance (Section 6.2), in which case the complexity w.r.t. LLM calls becomes $O(C_{max}\cdot N\cdot D_{max})$. Finally, compared to prior LLM-based document categorization methods (Wan et al., 2024; Pham et al., 2024), since AgenticTagger does not have sequential dependency between most of its dominant LLM operations (specifically, the $\pi_{\text{Annotator}}$ annotations), the overall execution time can be further reduced approximately linearly w.r.t. the number of concurrent threads in the ParallelAssign operation.
Table 1. AgenticTagger for Generative Recommendation on academic benchmarks. For each metric, the best result is in bold and the second-best is underlined. Our model’s performance is reported across 3 runs. We provide additional comparison against crossing-based IDs and free-form generations in Figure 3 and Section 6.3.
@@
Once the vocabulary is built, the assignment stage simply involves looping through all items of interest, and the assignment loop induces $O(|\mathcal{I}|)$ LLM calls, i.e., the cost of querying $\pi_{Annotator}$ in a single-pass manner over the item corpus. To this end, after the vocabulary is built, AgenticTagger can easily scale to large item corpora with parallelized calls to LLMs. Further, since the only difference across LLM calls during the assignment stage is the item description, the inference can, in principle, be further accelerated by infrastructure optimizations such as KV-Cache (Pope et al., 2023), which we leave for future work.
@@ 4.4. Modeling AgenticTagger Features
Once the process is complete, AgenticTagger assigns a fixed maximum number of descriptors to an item, where each “slot” represents a categorical feature. In this work, we opt to assume the best potential value for each slot is nominal, i.e., mutually exclusive, although there could be more than one suitable value at the same slot for each item (e.g., a song that is both “Rock” and “Blues”). These assigned descriptors also have an inherent order, where a preceding descriptor represents the broader, parent category of its child descriptors. For example, for an item to fit the “Glam Rock” descriptor, its preceding, broader feature would be “Rock”. As we later show, this formulation allows these features to be natively used as item IDs, such as semantic IDs and ID-like ranking features (Section 5).
@@ 5. Main Experiments
In this section, we showcase the performance of the AgenticTagger-produced descriptor features by deploying them as discrete ID features in ML models, in particular, generative recommendation (Section 5.1) and ranking (Section 5.2). We then discuss an interpretability-oriented application, namely critique-based generative recommendation (Section 5.3), before other ablation studies and analyses.
@@ 5.1. Generative Recommendation
We conduct our experiments across three domains from the Amazon Reviews dataset (McAuley et al., 2015): “Sports and Outdoor” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs); the dataset statistics are shown in Table 6. For the larger CDs domain only, we constrain AgenticTagger’s branching factor to 1 for scalability and apply the vanilla AgenticTagger algorithm to the other two domains (see Section 6.2 for discussions). We adopt the commonly used “last-out” splitting (Kang and McAuley, 2018; Rajput et al., 2023; Hou et al., 2025a), where we hold out the last chronologically interacted item for each user as the test set, and the second-last item as the validation set. Following recent work (Rajput et al., 2023; Hou et al., 2023; Wang et al., 2024a), we evaluate the performance of models with Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. To eliminate randomness in model optimization, we repeat our experiments with 3 random seeds and report the average metrics. We use Gemini-2.5-flash and Gemini-2.5-flash-lite for the architect and annotator LLMs, respectively, in our main experiment. We discuss results from Gemini-1.5-flash and Gemini-1.5-flash-lite (specifically, the 001 series), which achieved similarly competitive performance, in Section 6.4.
@@
In our main experiment (Table 1), we compare our method with several families of models: models with item IDs only, such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019); models enhanced by embedded features: FDSA (Zhang et al., 2019), $\text{S}^{3}$-Rec (Zhou et al., 2020); and models with discrete features (i.e., semantic IDs) derived from semantic information or collaborative information: VQ-Rec (Hou et al., 2023), TIGER (Rajput et al., 2023), HSTU (Zhai et al., 2024), and LLM-Indexer (Jin et al., 2024). We additionally compare our method with models that enhance raw semantic IDs via feature crossing, namely SPM-CID (Singh et al., 2024) and ActionPiece (Hou et al., 2025b) in Figure 3.
@@
Interestingly, the performance improvements are more significant in the Sports and CDs domains, with 8.9%-11.6% improvements over the closest competing model in NDCG@10, whereas the improvement is moderate (2%) in the Beauty domain. We hypothesize that this could be because AgenticTagger yields the most benefit when it is able to construct its vocabulary from a larger item corpus, where the natural language feedback aggregated via the Distill operation provides more stable signals. We further discuss this hypothesis in Section 6.2, where we show that AgenticTagger benefits from larger datasets with diverse items, even when doing so at the cost of pruning the vocabulary-building process.
@@
In light of recent advances in demonstrating the benefit of incorporating dataset collaborative filtering statistics directly into the tokenization process via feature crossing (Hou et al., 2025b; Singh et al., 2024), we also compare our method with two recent tokenization methods that merge frequently co-occurring semantic IDs into new IDs based on observed user behavior sequences, SentencePiece (Singh et al., 2024) and ActionPiece (Hou et al., 2025b). To be specific, these methods represent an alternative path towards better features for recommender systems: automated feature crossing based on user behavior. As shown in Figure 3, AgenticTagger remains a strong tokenization method, demonstrating the effectiveness of leveraging LLMs’ content understanding abilities in tokenization in addition to collaborative filtering signals. Meanwhile, we note that AgenticTagger is orthogonal to these crossing-based techniques, and we leave the crossing of AgenticTagger features to future study.
@@ 5.2. Ranking
Table 2. Ranking performance comparison between the Item ID baseline and AgenticTagger.
@@
R@5 0.2655 $\pm$ 0.0037 0.2823 $\pm$ 0.0038
N@5 0.1966 $\pm$ 0.0029 0.1870 $\pm$ 0.0027
R@10 0.3456 $\pm$ 0.0040 0.4300 $\pm$ 0.0041
N@10 0.2224 $\pm$ 0.0029 0.2346 $\pm$ 0.0027
R@20 0.4427 $\pm$ 0.0042 0.6082 $\pm$ 0.0041
N@20 0.2470 $\pm$ 0.0029 0.2796 $\pm$ 0.0025
R@50 0.5761 $\pm$ 0.0041 0.8514 $\pm$ 0.0030
N@50 0.2733 $\pm$ 0.0027 0.3280 $\pm$ 0.0022
5.2.1. Dataset, Baseline, and Evaluation.
Besides generative recommendation, another frequent use case for structured item features is ranking (Singh et al., 2024; Zhu et al., 2021). To this end, we evaluate the effect of AgenticTagger features on a private dataset collected from a news-feed serving platform over a random subset of 47,358 users across an 8-day period. This results in 67,097 items and 206,345 interactions, where raw item content and user interactions are fully accessible. Following prior work in the literature (He et al., 2017; Kang and McAuley, 2018), we randomly sample 100 negative items per user to reduce computational overhead, and report performance with Recall@$K$ and NDCG@$K$, where $K\in\{5,10,20,50\}$, similar to our experiments on public benchmarks. The backbone recommendation model is a neural feature-aware model aligned with the deployed model in the platform, where we use either one-hot item IDs (i.e., one embedding per item) or AgenticTagger descriptors as IDs during training. For good scalability, we adopt a 3-layer setting for AgenticTagger, and build the vocabulary using 30,000 randomly subsampled items, resulting in 3 descriptors per item. We break item collisions using an extra digit of a collision-resolver token following standard practice in the semantic ID literature (Rajput et al., 2023; Hou et al., 2025a; He et al., 2025; Singh et al., 2024).
5.2.2. Overall Performance.
We report our results in Table 2; as shown, AgenticTagger brings significant improvements to ranking accuracy compared to raw item IDs, showcasing the effectiveness of AgenticTagger features. On the other hand, an interesting observation is that the item ID baseline shows slightly stronger performance for NDCG@5. We hypothesize that this is because when the dataset is sufficiently large for good convergence, as in our case, memorization of item-specific collaborative filtering signals via unique per-item embeddings helps fine-grained ranking among top candidates. This is aligned with prior findings in the literature (Kang et al., 2021), where compressing one-hot item representations into hashed embedding tables sometimes requires (albeit marginally) a performance trade-off. We discuss further verification of this idea in the appendix, where we show that in smaller datasets where interactions are sparser, AgenticTagger-generated features exhibit larger improvements compared to one-hot item representations (Section B.2).
5.2.3. Cardinality.
As discussed in the previous paragraph, AgenticTagger achieves strong performance with a much lower-cardinality vocabulary, utilizing only 2.7k unique descriptors compared to the full one-hot item table of over 67k item embeddings, achieving effective vocabulary compression. This demonstrates that AgenticTagger effectively captures the distinctive characteristics of each item within its automatically inferred descriptor vocabulary, validating the effectiveness of our proposed framework in enabling LLMs to understand and capture frequent patterns in the item corpus.
5.3. Interpretability and Controllability
@@
As discussed earlier, a core advantage of LLM-generated features is their explainability, which naturally suits user-facing use cases, such as situations where one would like to control or constrain the matching mechanism of a recommender system. In this section, we discuss the value AgenticTagger features add to the interpretability and controllability of a recommender system via a single-turn critique setting, where a user gives feedback to the recommender system, providing information on the item they would like to interact with next. In particular, this setting is a fundamental feature in many conversational recommender systems (Li et al., 2018; Wang et al., 2022; He et al., 2023; Zhu et al., 2025b), and commonly manifests in production recommender systems as feedback with keyphrase-based rationales.
Inspired by recent work on using LLMs for user simulation in recommender systems (Bougie and Watanabe, 2025; Park, 2025; Zhang et al., 2025a; Wang et al., 2024f) and conversational recommendation (Wang et al., 2023; Zhu et al., 2025a), we make use of an LLM as a user simulator to evaluate whether AgenticTagger can enable effective user control over recommender system behavior. Specifically, we adopt a communication game in an information bottleneck setting, where an LLM (gemini-2.5-flash in practice) observes the descriptions of target items in its prompt, and is asked to select all potentially relevant top-level (i.e., level 1) features in the vocabulary, which generally contains a small number of keywords (e.g., 16 for the CDs and Vinyl domain, the largest public benchmark we evaluate on). We then generate the next item from a pre-trained generative recommender backbone using beam search, following standard practice, except we constrain its search space to be within the allowed level 1 features provided by the user simulator. While this interaction paradigm essentially “leaks” ground-truth information to the recommender, it verifies that (1) AgenticTagger produces a vocabulary that precisely identifies a meaningful subset of items, and (2) the produced vocabulary aligns well with the semantic understanding of LLMs, which are known to align well with human perception (Chang and Bergen, 2024), demonstrating the interpretability and controllability potential of AgenticTagger features.
As shown in Table 3, this constrained recommendation setting increases model performance significantly across datasets, showcasing that AgenticTagger features effectively communicate the nuances of items in the dataset and, importantly, in an interpretable manner, such that the behavior of the corresponding recommender system using AgenticTagger features can be straightforwardly controlled by natural-language-based critiques.